### v1

# Summary of the Precision Measurements of the Electroweak Mixing Angle in the Region of the Z pole

Abstract
[…]
1 Introduction
The effective leptonic weak mixing angle ($\sin^{2}\theta^{\ell}_{\mathrm{eff}}$) is a key parameter in precision tests of the Standard Model (SM). A recent CMS analysis of the forward-backward asymmetry ($A_{\rm FB}$) in Drell-Yan events at 13 TeV [26] achieved a landmark precision, comparable to results from electron-positron LEP/SLD colliders. Despite the high experimental precision achieved by the CMS analysis, uncertainties related to the proton’s parton distribution functions (PDFs) remain the dominant limitation.
This work summarizes the results obtained using an extended PDF profiling strategy applied to the CMS measurement. We incorporate additional, powerful constraints from CMS measurements of the $W$-boson decay lepton asymmetry [36] and $W/Z$ cross-section ratios [27]. After establishing consistency with the published CMS measurement, the extended fit strategy leads to a noticeable reduction in the PDF-driven component of the total uncertainty. This approach results in the most precise single measurement of $\sin^{2}\theta^{\ell}_{\mathrm{eff}}$ to date.
2 Reproduction of the CMS 13 TeV analysis
The CMS extraction of the weak mixing angle is based on a global fit framework that incorporates correlated experimental and theoretical uncertainties [26].
$\chi^{2}(\beta_{\rm exp},\beta_{\rm th})=\sum_{i}\frac{(\sigma_{i}^{\rm exp}+\sum_{j}\Gamma_{ij}^{\rm exp}\beta_{j,{\rm exp}}-\sigma_{i}^{\rm th}-\sum_{k}\Gamma_{ik}^{\rm th}\beta_{k,{\rm th}})^{2}}{\Delta_{i}^{2}}+\sum_{j}\beta^{2}_{j,{\rm exp}}+\sum_{k}\beta^{2}_{k,{\rm th}}.$ (1)
Figure 1: Extracted values of $\sin^{2}\theta^{\ell}_{\mathrm{eff}}$ from the 13 TeV CMS $A_{4}$ data for 19 different PDF sets on the horizontal axis. (A) Before profiling. (B) After profiling with $A_{4}$. (C) After profiling with $A_{4}$ plus $W$ decay lepton asymmetry. (D) After profiling with $A_{4}$ plus $W$ decay lepton asymmetry plus W/Z cross section ratios. The vertical axis shows the $\chi^{2}$ values of the fits divided by the number of degrees of freedom ($N_{\rm data}-1$), where the one degree of freedom corresponds to the free parameter $\sin^{2}\theta^{\ell}_{\mathrm{eff}}$.
Correlated experimental and theoretical uncertainties are implemented by introducing nuisance parameters, $\beta_{\rm exp}$ and $\beta_{\rm th}$, which propagate their effects to the measured spectra and theoretical predictions through the corresponding response matrices $\Gamma^{\rm exp}$ and $\Gamma^{\rm th}$. Within this framework, the index $i$ labels the $N_{\rm data}$ points of the $(|y|,M)$ double-differential $A_{4}$ dataset, corresponding to 63 entries used in the $A_{4}$ profiling. The indices $j$ and $k$ index the nuisance parameters associated with experimental and theoretical uncertainty sources, respectively. The measured values and their uncorrelated experimental uncertainties are denoted by $\sigma_{i}^{\rm exp}$ and $\Delta_{i}$, while the theoretical predictions are given by $\sigma_{i}^{\rm th}$. Correlations among experimental uncertainties in the double-differential $A_{4}$ measurement are accounted for by the covariance matrix $\Gamma^{\rm exp}$.
Modern QCD calculations are used to model the Drell–Yan process, following the setup of the original CMS analysis [31, 25, 22, 21, 15]. And the $Z$-boson transverse-momentum spectrum is reweighted to data, following the original CMS analysis.
The theoretical uncertainty accounts for effects arising from missing higher-order contributions in both QCD and electroweak (EW) calculations. Electroweak corrections are evaluated using POWHEG-Z_ew, while PDF-related uncertainties are derived from NLO grids produced with MadGraph5-aMC@nlo [8] and interfaced through PineAPPL [35, 20]. These components are incorporated into the matrix $\Gamma^{\rm th}$ via nuisance parameters that represent missing higher-order EW effects, PDF Hessian variations, and the effective weak mixing angle $\sin^{2}\theta^{\ell}_{\mathrm{eff}}$ itself, which is treated as a free parameter in the fit. Uncertainties associated with missing higher-order QCD corrections are assessed by repeating the fit under six independent scale variations, with the largest observed deviation from the nominal result taken as the corresponding uncertainty.
Table 1 summarizes the extracted values of $\sin^{2}\theta^{\ell}_{\mathrm{eff}}$, expressed in units of $10^{-5}$, obtained with the 13 TeV $A_{4}$ distribution both prior to and following the profiling procedure, for a total of 19 PDF sets. The majority of these PDFs are computed at next-to-next-to-leading-order (NNLO) accuracy in QCD, while a subset corresponds to approximate next-to-next-to-next-to-leading-order calculations (an3lo). PDF ensembles marked as “qed” incorporate Quantum Electrodynamics effects and explicitly include photon parton densities. Within the NNPDF40 family, the “mhou” designation denotes the inclusion of corrections associated with missing higher-order uncertainties. The MSHT and NNPDF sets adopt asymmetric parameterizations of the strange-quark sea, whereas the CTEQ sets assume a symmetric strange-quark sea, with the exception of the CT18As PDF.
The total uncertainties in Table 1 include contributions from statistical, experimental systematic, theoretical, and PDF sources. The differences in the results obtained with the various PDF sets (before profiling) largely reflect variations in input datasets, parameterization choices, and flavor assumptions. If there are sufficient parameters in the PDF sets, then the profiled PDFs should be able to describe all data that are used in the $\sin^{2}\theta^{\ell}_{\mathrm{eff}}$ profiling analysis with good $\chi^{2}$. The columns labeled marked "Diff. from CT18Z" show the difference between the value of $\sin^{2}\theta^{\ell}_{\mathrm{eff}}$ extracted with each of the 19 PDF sets and that obtained using the nominal CT18Znnlo PDF set.
For the nominal CT18Znnlo set, the profiled value obtained from the 13 TeV $A_{4}$ data alone is in very good agreement with the published CMS result (Table 1), which validates our implementation of the CMS analysis in xFitter [7, 19].
PDF $A_{4}$ without profiling $A_{4}$ profiling $A_{4}+W_{\rm asym}+W/Z$ profiling
(63 data points) (63 data points) (83 data points)
$\sin^{2}\theta^{\ell}_{\mathrm{eff}}$ Diff. $\chi^{2}$ $\sin^{2}\theta^{\ell}_{\mathrm{eff}}$ Diff. $\chi^{2}$ $\sin^{2}\theta^{\ell}_{\mathrm{eff}}$ Diff. $\chi^{2}$
from from from
CT18Z CT18Z CT18Z
NNPDF40
nnlo_as_01180_hessian [10] 23107 $\pm$ 49 -59 89 23130 $\pm$ 24 -22 66 23152 $\pm$ 23 -4 83
nnlo_as_01180_qed [13] 23102 $\pm$ 46 -64 108 23144 $\pm$ 23 -8 69 23161 $\pm$ 22 5 87
nnlo_as_01180_mhou [12] 23101 $\pm$ 45 -65 79 23110 $\pm$ 27 -42 64 23145 $\pm$ 24 -11 88
nnlo_as_01180_qed_mhou [13] 23103 $\pm$ 46 -63 93 23136 $\pm$ 24 -16 69 23156 $\pm$ 23 0 88
an3lo_as_01180 [14] 23129 $\pm$ 46 -37 72 23132 $\pm$ 25 -20 65 23154 $\pm$ 23 -2 87
an3lo_as_01180_qed [14] 23129 $\pm$ 49 -37 78 23147 $\pm$ 23 -5 68 23160 $\pm$ 22 4 89
an3lo_as_01180_mhou [14] 23106 $\pm$ 47 -60 78 23120 $\pm$ 26 -32 65 23140 $\pm$ 25 -16 86
an3lo_as_01180_qed_mhou [14] 23102 $\pm$ 44 -64 90 23137 $\pm$ 23 -15 69 23156 $\pm$ 23 0 91
CTEQ
CT18nnlo [29] 23204 $\pm$ 64 38 92 23166 $\pm$ 36 14 64 23169 $\pm$ 24 13 84
CT18Znnlo [29] 23166 $\pm$ 56 0 70 23152 $\pm$ 32 0 65 23156 $\pm$ 24 0 79
CT18Annlo [29] 23179 $\pm$ 54 13 76 23164 $\pm$ 28 12 67 23160 $\pm$ 23 4 79
CT18Xnnlo [29] 23194 $\pm$ 53 28 86 23171 $\pm$ 30 19 64 23160 $\pm$ 24 4 87
CT18qed-proton [38] 23204 $\pm$ 64 38 93 23164 $\pm$ 36 12 64 23155 $\pm$ 24 -1 80
CT18As_LatNNLO [28] 23177 $\pm$ 75 11 74 23108 $\pm$ 43 -44 64 23147 $\pm$ 35 -9 80
MSHT20
MSHT20nnlo_as118 [9] 23053 $\pm$ 46 -113 200 23115 $\pm$ 30 -37 75 23123 $\pm$ 26 -33 96
MSHT20qed_nnlo [23] 23055 $\pm$ 59 -111 188 23124 $\pm$ 31 -28 73 23145 $\pm$ 27 -11 90
MSHT20an3lo_as118 [30] 23101 $\pm$ 47 -65 138 23138 $\pm$ 29 -14 69 23154 $\pm$ 26 -2 79
MSHT20qed_an3lo [24] 23101 $\pm$ 52 -65 126 23137 $\pm$ 31 -15 69 23143 $\pm$ 28 -13 80
PDF4LHC21_40 [11] 23122 $\pm$ 83 -44 80 23133 $\pm$ 33 -19 62 23143 $\pm$ 27 -13 77
Table 1: Values of $\sin^{2}\theta^{\ell}_{\mathrm{eff}}$ (in units of $10^{-5}$) before and after profiling with the CMS 13 TeV $A_{4}$ distribution (63 points) for 19 PDF sets. Also shown are the values extracted by including the CMS 13 TeV $W$-decay lepton asymmetry and the CMS W/Z cross section ratios (total of 83 points) in the profiling. The total uncertainties include contributions from statistical, experimental systematic, theoretical and PDF sources. The column marked "Diff. from CT18Z" is the difference from $\sin^{2}\theta^{\ell}_{\mathrm{eff}}$ extracted with the CT18Znnlo PDF set. Note, the $\chi^{2}$ values before profiling do not include PDF errors, while the $\chi^{2}$ values after profiling include PDF errors.
3 Extended Profiling with $W$ Charge Asymmetry and $W/Z$ Cross Section Ratios
After profiling, the dominant PDF uncertainty is largely reduced, but small differences between modern PDF sets are still observed. As shown in [17], including the $W$-boson charge asymmetry measurements ($W_{\rm asym}$) in the profiling provides additional constraints on the $d/u$ ratio, leading to a further reduction of the PDF uncertainty in $\sin^{2}\theta^{\ell}_{\mathrm{eff}}$. Moreover, some CT18 PDF sets lack sufficient constraints on the strange-quark distribution at high energy scales, resulting in systematically lower strange-quark densities compared to other PDFs and thereby predicting a larger effective weak mixing angle. Incorporating the $W/Z$ cross section ratios can provide further constraints on the strange-quark distribution, helping to reduce these discrepancies.
Therefore, we perform an updated analysis of the published CMS 13 TeV $A_{4}$ data (63 data points). We also include the recent CMS $W_{\rm asym}$ measurement [36] at 13 TeV (18 additional data points) and the CMS measurements of the $W$ and $Z$ fiducial cross-section ratios [27] at 5.02 TeV ($12.505\pm 0.037_{\rm stat}\pm 0.032_{\rm syst}$) and 13 TeV ($12.078\pm 0.028_{\rm stat}\pm 0.032_{\rm syst}$), contributing 2 additional data points. The results are shown in the columns of Table 1 and in Figure 1.
Extracted values of $\sin^{2}\theta^{\ell}_{\mathrm{eff}}$ from the 13 TeV CMS $A_{4}$ data for different PDF sets are shown on the horizontal axis of each of the four panels in Fig. 1. The vertical axis shows the values of $\chi^{2}$ normalized to the number of degrees of freedom, $\rm{ndf}=N_{\rm data}-1$, where the subtraction accounts for the single free parameter $\sin^{2}\theta^{\ell}_{\mathrm{eff}}$. Shown are the values before profiling (panel A, 63 data points). After profiling with $A_{4}$ (panel B, 63 data points), after profiling with $A_{4}$ and also with $W$ decay lepton asymmetry (panel C, 81 data points), and after profiling with $A_{4}$ and the $W$ decay lepton asymmetry and also the W/Z cross section ratios (panel D, 83 data points).
As seen in Table 1 and Figure 1, after $A_{4}$ profiling the error in the extracted value of $\sin^{2}\theta^{\ell}_{\mathrm{eff}}$ with the nominal CT18Znnlo PDF set is reduced from 0.00056 to 0.00032. By also including the $W$ decay lepton asymmetry and the $W/Z$ cross section ratios the uncertainty for the nominal CT18Znnlo is reduced to 0.00024.
Figure 2: Comparison of $\sin^{2}\theta^{\ell}_{\mathrm{eff}}$ extracted in this analysis (labeled "This analysis 13 TeV") with previous measurements [34, 5, 6, 1, 37, 26, 3, 2] and the prediction of the 2025 SM global fit [32, 33]. Also shown is the prediction of the Two Higgs Doublet Model [16] corresponding to the CDF $M_{W}$ value [4] (80.4335 $\pm$0.0094 GeV).
Including all three inputs in the profiling procedure results in a pronounced alignment among the different PDF sets. Following profiling, the extracted central values from 18 of the 19 PDFs (including three of the four MSHT20 sets) are consistent with the CT18Znnlo determination at the level of one standard deviation. The only exception is the profiled MSHT20nnlo_as118 set, whose central value lies 1.38 standard deviations below the CT18Znnlo result and is associated with a poor $\chi^{2}$. By contrast, as reported in Table 1, the profiled MSHT20an3lo_as118 set yields an identical value of $\sin^{2}\theta^{\ell}_{\mathrm{eff}}$ to that obtained with CT18Znnlo, together with an acceptable $\chi^{2}$.
Figure 2 shows a comparison of $\sin^{2}\theta^{\ell}_{\mathrm{eff}}$ extracted in this analysis with the nominal CT18Znnlo PDF set (Labeled "This analysis 13 TeV") to previous measurements [34, 5, 6, 1, 37, 26, 3, 2]. Also shown are the predictions of the 2025 SM global fit [32, 33] and the prediction of the Two Higgs Doublet Model [16] assuming the CDF $M_{W}$ value [4] of 80.4335 $\pm$0.0094 GeV.
4 Conclusion
[…]
Research supported by the U.S. Department of Energy under University of Rochester grant number DE-SC0008475 and by National Research Foundation of Korea (NRF) grants (RS-2024-00350406, RS-2008-NR007227).

### v2

@@
 4 Conclusion
-[…]
+This proceeding reports on an improved extraction of the effective leptonic weak mixing angle, $\sin^{2}\theta^{\ell}_{\mathrm{eff}}$, in which additional CMS measurements of the $W$-boson charge asymmetry and $W/Z$ cross-section ratios are incorporated to further constrain parton distribution function effects.
+The final result, $\sin^{2}\theta^{\ell}_{\mathrm{eff}}=0.23156\pm 0.00024$, represents the most precise single measurement of this parameter to date, as presented in Ref.
+[18].
+A key outcome of this enhanced PDF profiling is the excellent consistency observed across 18 different PDF sets, with all results lying within one standard deviation of the nominal CT18ZNNLO value.
+The extracted value is compatible with the Standard Model expectation, $\sin^{2}\theta^{\ell}_{\mathrm{eff}}=0.23161\pm 0.00004$.
 Research supported by the U.S.